OpenStudy (anonymous):

solve the absolute value equation |x-3|-5=-3

OpenStudy (tkhunny):

The really important observation is this: \(|x-3| = (x-3)\) wherever \(x-3 \ge 0\) or \(x \ge 3\) \(|x-3| = -(x-3) = (3-x)\) wherever \(x < 3\) Now you have two simpler problems to work on.

OpenStudy (anonymous):

And what is those simpler problems?

OpenStudy (tkhunny):

You will need to use the substitution principle and the ideas just presented. Here is your problem statement: Solve: |x-3|-5=-3 If you PROMISE \(x \ge 3\), this is your new problem statement" Solve: (x-3)-5=-3 Can you solve this one?

OpenStudy (anonymous):

no :/ i don't.

OpenStudy (tkhunny):

Oh, well then it seems you are in the wrong class or you have been sleeping in class. Let's start from the beginning. Can you solve this: x - 3 = 2

OpenStudy (anonymous):

x=5

OpenStudy (tkhunny):

Perfect. How about this one? 4x - 12 = 16

OpenStudy (anonymous):

x=7

OpenStudy (tkhunny):

One more: (x-3)-5=-3

OpenStudy (anonymous):

x=5

OpenStudy (tkhunny):

We are now ready to solve your problem: If you PROMISE x≥3 , this is your new problem statement: Solve: (x-3)-5=-3 Can you solve this one?

OpenStudy (anonymous):

x=5 right?

OpenStudy (tkhunny):

Is 5 greater than 3? If so, then you have kept your promise and you have a solution. Now this one. If you PROMISE x<3 , this is your new problem statement: Solve: (3-x)-5=-3

OpenStudy (anonymous):

x=1

OpenStudy (tkhunny):

(3-x) - 5 = -3 Add 5 (3-x) = 2 Add x 3 = 2+x Subtract 2 1 = x Good. Is 1 less than 3? If so, you have kept your promise and you have another solution. Now, go try 5 and 1 in the original problem statement and make sure they work!

OpenStudy (anonymous):

thnx i already have my answer